# Instant Runoff Voting and the Reinforcement Paradox

### v2

@@ 3.1. Consequences of the Algebraic Conditions
Consider the profile in Table 6 where $A$ is the IRV winner. The election demonstrates a runner-up paradox and $B$ wins in both sub-elections $\mathcal{B}_{1}$ and $\mathcal{B}_{2}$. However each of these sub-elections themselves satisfies the conditions for a runner-up paradox. This means that sub-election $\mathcal{B}_{1}$, in which $C$ is the plurality loser and $A$ is the runner-up, can be further subdivided into two smaller sub-elections in which $A$ is the IRV winner. (One such partition is $\mathcal{B}_{11}$ which corresponds to $X_{1}=12$, $X_{2}=31$, $Y_{1}=0$, $Y_{2}=25$, $Z_{1}=0$ and $Z_{2}=17$ and $\mathcal{B}_{12}$ which corresponds to $X_{1}=28$, $X_{2}=31$, $Y_{1}=5$,$Y_{2}=26$, $Z_{1}=1$ and $Z_{2}=31$.) And sub-election $\mathcal{B}_{2}$, in which $A$ is the plurality loser and $C$ is the runner-up, can be further subdivided into two smaller sub-elections in which $C$ is the IRV winner. (One such partition is $\mathcal{B}_{21}$ which corresponds to $X_{1}=50$, $X_{2}=0$, $Y_{1}=4$, $Y_{2}=25$, $Z_{1}=30$ and $Z_{2}=0$ and $\mathcal{B}_{22}$ which corresponds to $X_{1}=10$, $X_{2}=38$, $Y_{1}=41$, $Y_{2}=54$, $Z_{1}=69$ and $Z_{2}=2$.)
@@
1st choice $A$ $A$ $B$ $B$ $C$ $C$
2nd choice $B$ $C$ $A$ $C$ $A$ $B$
3rd choice $C$ $B$ $C$ $A$ $B$ $A$
@@
1st choice $A$ $B$ $C$ $C$
2nd choice $C$ $C$ $A$ $B$
3rd choice $B$ $A$ $B$ $A$
@@
Finally, we note that elections may be susceptible to a plurality loser paradox but not a runn-erup paradox, as demonstrated by the profile in Table 7 where neither Condition (1) nor (2) of Proposition 1 is satisfied, but the plurality loser inequalities are satisfied. Such preference profiles can arise when voter preferences are “single-peaked.” In this case, candidate $C$ is like a “centrist” candidate in a spatial model, as if $A$, $B$, and $C$ have been arranged on a left-to-right line, with $C$ in the middle.
Of course, real-world elections typically do not return ballot data resembling the profile in Table 7. In fact as we will see in Section 6 in real elections, the IRV winner tends to be a much stronger candidate than the plurality loser. Thus we expect that if an election demonstrates a plurality loser reinforcement paradox in this data then the election will also demonstrate a runner-up paradox.
@@ 4. Threshold Results
In this section we examine the minimal percentage each candidate can earn while the election still demonstrates a reinforcement paradox. For some classical paradoxes such as upward and downward monotonicity paradoxes, an election can exhibit such a paradox only if each of the three candidates is not “weak.” We show this is not the case for a general reinforcement paradox in Proposition 3, where the plurality loser can have an arbitrarily low percentage of the first-places votes and yet an election can demonstrate a reinforcement paradox. We show in Proposition 4 that the candidate affected by the paradox must earn at least 25% of the first place vote, so that if an election demonstrates a runner-up (respectively plurality loser) reinforcement paradox then the IRV runner-up (respectively plurality loser) must earn at least 25% of the first-place votes.
In this section we assume that the number of voters approaches infinity, so that we use proportions of votes rather than integers. We also assume that each voter provides a complete ranking as this does not affect the logic of the proofs.
@@
The plurality loser can have arbitrarily small first-place support and yet the election demonstrate the reinforcement paradox.
@@
Consider the profile in Table 8, where $p_{A}$ is the proportion of voters who cast the ballot $A\succ B\succ C$ and also the proportion of voters who cast the ballot $A\succ C\succ B$, and $\epsilon$ and $\delta$ are positive and very small. Furthermore, assume $0.52p_{A}+\epsilon<$ 0.5 and $14p_{A}+\epsilon+\delta=$ 1 (which implies $\delta>\epsilon)$. Note that $A$ is the IRV winner and $C$ is the plurality loser.
This profile satisfies condition (2) of Proposition 1 and thus the election demonstrates a runner-up reinforcement paradox. For completeness, we include a partition of the electorate into two sub-electorates, as shown in the table. Let $\delta_{B}$ and $\delta_{A}$ be positive small parameters such that $20<\delta_{B}<\delta_{A}<\delta/$ 2. Then $B$ wins when using the ballots of both sub-electorates separately. Since $\delta$ is arbitrary, the proposition is proven.∎
Prop. Voters $p_{A}$ $p_{A}$ $2p_{A}+\epsilon$ $\delta$
1st choice $A$ $A$ $B$ $C$
2nd choice $B$ $C$ $A$ $A$
@@
$\mathcal{B}_{2}$ $p_{A}-\delta+\delta_{A}$ $p_{A}$ $2p_{A}+\epsilon-\delta+\delta_{B}$ 0
@@
Thus, for a large enough electorate the plurality loser can control an arbitrarily small amount of first-place votes and the election can demonstrate a reinforcement paradox. This proposition distinguishes reinforcement paradoxes from other classical paradoxes such as upward or downward monotonicity paradoxes. For a three-candidate election to demonstrate an upward (resp. downward) paradox, each candidate must earn at least 25% (respectively $16\frac{2}{3}\$) of the first-place votes [29]. That is, for some classical paradoxes none of the candidates can have an extremely weak first-place showing; each candidate must be “viable” in some sense. Proposition 3 shows this is not the case for the reinforcement paradox. However, in the next proposition we show that for a runner-up (resp. plurality) reinforcement paradox, the runner-up (resp. plurality loser) must earn at least 25% of the first-place votes.
@@
The IRV runner-up (resp. plurality loser) can have, in the limit, as little as 25% of the first-place vote and the election demonstrate a runner-up (resp. plurality loser) reinforcement paradox. In addition, the IRV winner can have, in the limit, as much as 50% of the first-place vote and the election demonstrate either a runner-up or plurality loser reinforcement paradox.
@@
To see that that bounds are tight, consider the profile in Table 9 where $\epsilon_{A},\epsilon_{B},\epsilon_{1}$ and $\epsilon_{2}>0$ are arbitrarily small parameters. If $\epsilon_{B}>\epsilon_{2}-\epsilon_{A}$ then this profile satisfies the conditions in (2) of Proposition 1.
@@
1st choice $A$ $B$ $C$ $C$
2nd choice $B$ $A$ $A$ $B$
3rd choice $C$ $C$ $B$ $A$
@@
1st choice $A$ $B$ $C$
2nd choice $C$ $C$ $A$
3rd choice $B$ $A$ $B$
@@
We conclude this section by noting that there are elections that demonstrate a reinforcement paradox in which one of the sub-elections is arbitrarily small.
@@
Let $m\in\mathbb{N}$ be a large even number. Consider an election with ballot set $\mathcal{B}$ where $m$ ballots have the form $A\succ B$, $m+1$ have the form $B\succ A$, $m/2$ have the form $C\succ A$, and $m/2-2$ have the form $C\succ B$. Then $A$ wins this election, since initially they have one fewer vote than $B$ but the elimination of $C$ allows $A$ to win by one vote.
Let $\mathcal{B}_{1}$ consist of two ballots of the form $A\succ B$, three of the form $B\succ A$, and four of the form $C\succ A$. Let $\mathcal{B}_{2}$ consist of $m-2$ ballots of the form $A\succ B$, $m-2$ of the form $B\succ A$, $m/2-4$ of the form $C\succ A$, and $m/2-2$ of the form $C\succ B$. Then $\mathcal{B}_{1}$ and $\mathcal{B}_{2}$ form a partition of $\mathcal{B}$, and candidate $B$ wins both of these sub-elections. Since $m$ can be arbitrarily large, this implies there exists an election demonstrating a reinforcement paradox where one of the sub-elections can be arbitrarily small.
@@ 5. Results under various models
In this section we use different models of voter behavior to generate millions of synthetic elections. Among the many possible models we could use, we choose models which have been used extensively in prior social choice research. For each model, we generate 100,000 profiles and check whether a generated election demonstrates a reinforcement paradox. Thus, our primary tool for providing (estimated) probabilities is Monte Carlo simulation, although we can provide exact probabilities using other means for the impartial anonymous culture (IAC) model.
We are interested in probabilities for large electorates, as the real-world elections we analyze in the next section generally have thousands of voters. Thus, our simulations use an electorate size of $V=1001$ voters. We also ran simulations with $V\in\{601,2003,3901\}$, but the estimated probabilities were not markedly different for these other choices of $V$. With the exception of the IAC model, in this section we consider probabilities only for the complete ballot case.
@@ 5.1. Descriptions of the Models
We use two different types of models: one-dimensional spatial models and models which sample over the ballot simplex. In the spatial framework, we consider three different models. The first model, which we denote 1D Spatial, places three candidates and $V=1001$ voters on the real line using the standard normal distribution, and a voter’s preference ranking is determined by their Euclidean distance from each candidate. For example, if the three candidates $A$, $B$, and $C$ are positioned by the normal distribution at $-1.2$, $-0.1$, and 1.1 respectively, then a voter positioned at 0.3 would cast a ballot with $B$ ranked first, $C$ ranked second, and $A$ ranked third.
Because of the increased political polarization in the United States and elsewhere, we also use two bimodal spatial models to investigate the potential effects of polarization on the probability of the reinforcement paradox. Under the second spatial model, which we denote 1D Bimodal, we create a bimodal distribution from an equally weighted combination of the normal distribution centered at $-2$ and the normal distribution centered at 2, both with standard deviation 1. See Figure 1(top) for a visualization of this bimodal distribution. We then place candidates and voters along the line using the generated (normalized) bimodal distribution as a probability distribution, similar to 1D Spatial. Under the third spatial model, which we denote 1D Bimodal Weighted, we use the same two normal distributions to generate the bimodal distribution but give 50% more weight to the left distribution, resulting in a bimodal distribution as shown Figure 1(bottom). We use this model to capture polarization when one side of the political spectrum is stronger than the other.
@@
For the ballot simplex models we use Dirichlet distributions with parameter $\alpha$, denoted Dir$(\alpha)$, and we use the implementation of these models provided by the software package VoteKit [10]. In a ballot simplex model, we place a probability distribution on the discrete ballot 5-simplex, which is the set of 6-tuples $(X_{1},X_{2},Y_{1},Y_{2},Z_{1},Z_{2})$ such that each number is non-negative and the six numbers sum to $V$. Under a Dirichlet model with $\alpha\in(0,\infty)$, as $\alpha\rightarrow 0$ the generated profile tends to weigh heavily in favor of a single candidate, so that the resulting election is a landslide victory for one of the candidates. For example, when $\alpha$ is close to zero, if $V=1001$ then a 6-tuple $(X_{1},X_{2},Y_{1},Y_{2},Z_{1},Z_{2})$ where $X_{1}+X_{2}>800$ is much more likely to be generated than a 6-tuple where each of the six numbers is approximately equal. When $\alpha=1$ the model places a uniform distribution on the simplex, so that any 6-tuple of ranking numbers is as likely as any other. Dir$(1)$ is referred to as the IAC model in social choice. As $\alpha$ increases the sampled 6-tuples tend to cluster in the center of the simplex, so that it becomes increasingly unlikely that a candidate earns a majority of first-place votes. In the limit $\alpha\rightarrow\infty$, the model is very similar to what is called the impartial culture (IC) model. Under the IC model, each voter independently chooses one of the six preference rankings at random, leading to elections that, with high probability, are closely tied. Thus, as $\alpha$ increases a typical generated election becomes highly competitive between the three candidates.
@@ 5.2. Results
We begin by providing exact probabilities under Dir$(1)$, the IAC model. This model is important in the social choice literature and therefore many researchers have dedicated significant effort to developing tools for calculating probabilities under this model. One of the fruits of these labors is the software package Normaliz [5], which can provide limiting probabilities under IAC in three-candidate elections given a set of inequalities (such as those given in Propositions 1 and 2). Normaliz can be used in the complete ballot case, where profiles $(X_{1},X_{2},Y_{1},Y_{2},Z_{1},Z_{2})$ are chosen at random from the 5-simplex, and in the partial ballot case, where profiles $(X_{1},X_{2},X_{3},Y_{1},Y_{2},Y_{3},Z_{1},Z_{2},Z_{3})$ are chosen at random from the 8-simplex. The probabilities in the following proposition were obtained directly from Normaliz by imputting the appropriate sets of inequalities.
@@
The limiting probabilities of a reinforcement paradox occurring under the Dir$(1)$ model (the IAC model) are as shown in Table LABEL:IAC_results for complete and partial ballots. The “Cond. Prob.” probabilities are conditioned on the absence of a majority candidate.
@@
Table 11. Probabilities under the Dir$(1)$ (IAC) model. The top two rows correspond to the complete ballot case; the bottom two rows correspond to the partial ballot case.
The probabilities of one or more paradoxes occurring are similar in the complete and partial ballots cases. When conditioned on the absence of a majority candidate, the probabilities are significantly lower in the partial ballot case. This is because the probability an election does not contain a majority candidate is $145/256=56.64\%$ in the 8-simplex but is only $7/16=43.75\%$ in the 5-simplex.
For finite $\alpha\neq 1$ we are unaware of analytical techniques to calculate exact probabilities, and therefore we use Monte Carlo simulation. Figure 2 illustrates the probabilities obtained from simulations for Dir$(\alpha)$ where $\alpha\in\{0.1,0.2,0.3,\dots,4.0\}$: the top image indicates unconditioned probabilities and the bottom image indicates probabilities conditioned on the absence of a majority candidate. The blue squares give the probability that an election demonstrates a reinforcement paradox (of either type), the red disks give the probability an election demonstrates a runner-up paradox, the black diamonds give the probability an election demonstrates a plurality loser paradox, and the green pluses give the probability an election demonstrates a double paradox.
Figure 2. Estimated probabilities that an election demonstrates a reinforcement paradox under the model Dir$(\alpha)$. The top figure shows unconditioned probabilities while the bottom figure shows probabilities conditioned on the absence of a majority candidate.
@@
Interestingly, the corresponding probabilities under 1D Spatial and 1D Bimodal are approximately equal, perhaps because both models are symmetric about zero. When we unbalance the distribution under 1D Bimodal Weighted, the probabilities change significantly. In particular, 1D Bimodal Weighted is the only model in this study under which it is more likely to observe a plurality loser paradox than a runner-up paradox. We do not know with certainty the reason for this, but we speculate that it is due to different responses to the so-called “center squeeze” [39]. In one-dimensional three-candidate spatial models, the candidate lying in the center will often be the plurality loser even if they are closest to the “median voter.” Using the notation of Section 3, if $C$ is the center candidate then $X_{1}=Y+1=0$ since these voters will second-rank the nedian candidate. Hence the conditions in Case (1) of Proposition (1) cannot be satisfied (since they imply $B$ has the majority of the first-place vote). The conditions in Case (2) reduce to
@@
If the distribution of voters is symmetric, $Z_{1}$ and $Z_{2}$ are likely to be similarly-sized, and so the rightmost inequality will be satisfied whenever $Y>X$. Thus a runner-up (RU) paradox will occur often when the IRV and plurality winners are distinct.
If the distribution of voters is not symmetric, as with the 1D Bimodal Weighted model, the left and right-wing candidates are likely to be near the left and right peaks of the distribution respectively, with the centrist in between but on the “upward slope” of the larger peak (the left one in Figure 1). Thus, the quantity $Z_{1}-Z_{2}$ is more likely to be larger than in the symmetric case, decreasing the likelihood of Case (2) occurring.
A similar argument holds if $A$ or $B$ is the centrist candidate.
For the plurality loser (PL) paradox to occur, it is easily seen that $C$ must be the centrist candidate. In this case, $C$ is again likely to be on the upward slope of the larger peak in the 1D Bimodal Weighted model (and hence $Z$ is likely to be larger) than in the 1D Bimdodal model. So the probability of the plurality loser (PL) paradox occurring is greater in the nonsymmetric case than the symmetric.

### v3

@@ 2.1. Consequences of the Algebraic Conditions
Consider the profile in Table 6 where $A$ is the IRV winner. The election demonstrates a runner-up paradox and $B$ wins in both sub-elections $\mathcal{B}_{1}$ and $\mathcal{B}_{2}$. However each of these sub-elections themselves satisfies the conditions for a runner-up paradox. This means that sub-election $\mathcal{B}_{1}$, in which $C$ is the plurality loser and $A$ is the runner-up, can be further subdivided into two smaller sub-elections in which $A$ is the IRV winner. (One such partition is $\mathcal{B}_{11}$ which corresponds to $X_{1}=12$, $X_{2}=31$, $Y_{1}=0$, $Y_{2}=25$, $Z_{1}=0$ and $Z_{2}=17$ and $\mathcal{B}_{12}$ which corresponds to $X_{1}=28$, $X_{2}=31$, $Y_{1}=5$$Y_{2}=26$, $Z_{1}=1$ and $Z_{2}=31$.) And sub-election $\mathcal{B}_{2}$, in which $A$ is the plurality loser and $C$ is the runner-up, can be further subdivided into two smaller sub-elections in which $C$ is the IRV winner. (One such partition is $\mathcal{B}_{21}$ which corresponds to $X_{1}=50$, $X_{2}=0$, $Y_{1}=4$, $Y_{2}=25$, $Z_{1}=30$ and $Z_{2}=0$ and $\mathcal{B}_{22}$ which corresponds to $X_{1}=10$, $X_{2}=38$, $Y_{1}=41$, $Y_{2}=54$, $Z_{1}=69$ and $Z_{2}=2$.)
@@
1st choice AA AA $B$ $B$ $C$ $C$
2nd choice $B$ $C$ AA $C$ AA $B$
3rd choice $C$ $B$ $C$ AA $B$ AA
@@
1st choice AA $B$ $C$ $C$
2nd choice $C$ $C$ AA $B$
3rd choice $B$ AA $B$ AA
@@
Finally, we note that elections may be susceptible to a plurality loser paradox but not a runner-up paradox, as demonstrated by the profile in Table 7 where neither Condition (1) nor (2) of Proposition 1 is satisfied, but the plurality loser inequalities are satisfied. Such preference profiles can arise when voter preferences are “single-peaked.” In this case, candidate $C$ is like a “centrist” candidate in a spatial model, as if $A$, $B$, and $C$ have been arranged on a left-to-right line, with $C$ in the middle.
Of course, real-world elections typically do not return ballot data resembling the profile in Table 7. In fact as we will see in Section 5 in real elections, the IRV winner tends to be a much stronger candidate than the plurality loser. Thus, we expect that if an election is susceptible to a plurality loser reinforcement paradox in this data then it will also be susceptible to a runner-up paradox.
@@ 3. Threshold Results
In this section we examine the minimal percentage each candidate can earn while the election still demonstrates a reinforcement paradox. For some classical paradoxes such as upward and downward monotonicity paradoxes, an election can exhibit such a paradox only if each of the three candidates is not “weak.” We show this is not the case for a general reinforcement paradox in Proposition 3, where the plurality loser can have an arbitrarily low percentage of the first-places votes and yet an election is susceptible to a (runner-up) reinforcement paradox. We show in Proposition 4 that for an election to be susceptible to a runner-up (respectively plurality loser) paradox, the IRV runner-up (respectively plurality loser) must earn at least 25% of the first-place votes.
In this section, we assume that the number of voters approaches infinity, so that we use proportions of votes rather than integers. We also assume that each voter provides a complete ranking as this does not affect the logic of the proofs. Since the number of voters approaches infinity, we ignore the “+1” which appears in the Propositions in Section 2, as this term limits to zero.
@@
For any $\epsilon>0$, there is an election which is susceptible to the reinforcement paradox and the first-place vote share for the plurality loser is less than $\epsilon$.
@@
Consider the profile in Table 8, where $p_{A}$ is the proportion of voters who cast the ballot $A\succ B\succ C$ and also the proportion of voters who cast the ballot $A\succ C\succ B$, and $\beta$ and $\delta$ are positive and very small. Furthermore, assume $2p_{A}+\beta<0.5$ and $4p_{A}+\beta+\delta=1$ (which implies $\delta>\beta)$. Note that $A$ is the IRV winner and $C$ is the plurality loser.
This profile satisfies condition (2) of Proposition 1 and thus the election demonstrates a runner-up reinforcement paradox. Since $\delta$ is arbitrary, we can make $\delta<\epsilon$ and the proposition is proven. For completeness, we include a partition of the electorate into two sub-electorates, as shown in the table. Let $\delta_{B}$ and $\delta_{A}$ be positive small parameters such that $0<\delta_{B}<\delta_{A}<\delta/2$. Then $B$ wins when using the ballots of both sub-electorates separately. ∎
Prop. Voters $p_{A}$ $p_{A}$ $2p_{A}+\beta$ $\delta$
1st choice AA AA $B$ $C$
2nd choice $B$ $C$ AA AA
@@
$\mathcal{B}_{2}$ $p_{A}-\delta+\delta_{A}$ $p_{A}$ $2p_{A}+\beta-\delta+\delta_{B}$ 0
@@
Thus, for a large enough electorate the plurality loser can control an arbitrarily small amount of first-place votes and the election can demonstrate a reinforcement paradox. This proposition distinguishes reinforcement paradoxes from other classical paradoxes such as upward or downward monotonicity paradoxes. For a three-candidate election to demonstrate an upward (resp. downward) paradox, each candidate must earn at least 25% (respectively $16\frac{2}{3}\%$) of the first-place votes [39]. That is, for some classical paradoxes none of the candidates can have an extremely weak first-place showing; each candidate must be “viable” in some sense. Proposition 3 shows this is not the case for the reinforcement paradox. However, in the next proposition we show that for a runner-up (resp. plurality) reinforcement paradox, the runner-up (resp. plurality loser) must earn at least 25% of the first-place votes.
@@
If an election is susceptible to a runner-up (resp. plurality loser) reinforcement paradox then the runner-up (resp. plurality loser) must earn at least 25% of the first-place votes in round 1. This bound is tight in the limit as the number of voters approaches $\infty$.
@@
To see that that bounds are tight, consider the profile in Table 9 where $\epsilon_{A},\epsilon_{B},\epsilon_{1}$ and $\epsilon_{2}\geq 0$ are small parameters satisfying $\epsilon_{B}=\epsilon_{A}-\epsilon_{1}+\epsilon_{2}.$ If $\epsilon_{B}>\epsilon_{2}-\epsilon_{A}$ then this profile satisfies the conditions in (2) of Proposition 1.
@@
1st choice AA $B$ $C$ $C$
2nd choice $B$ AA AA $B$
3rd choice $C$ $C$ $B$ AA
@@
1st choice AA $B$ $C$
2nd choice $C$ $C$ AA
3rd choice $B$ AA $B$
@@
We conclude this section by noting that there are elections susceptible to a reinforcement paradox in which one of the sub-elections has an arbitrarily small percentage of voters.
@@
Let $m\in\mathbb{N}$ be a large even number. Consider an election $\mathcal{B}$ with 3m13m-1 ballots (complete or partial) in which the first and second preferences are distributed as follows: $m$ ballots of the form $A\succ B$; $m+1$ ballots of the form $B\succ A$; $m/2$ ballots of the form $C\succ A$; and m/22m/2-2 ballots of the form $C\succ B$. Candidate $A$ wins this election in two rounds after the $C$ is eliminated in round 1.
Let $\mathcal{B}_{1}$ consist of the sub-election with $9$ ballots: $2$ ballots of the form $A\succ B$; $3$ ballots of the form $B\succ A$; and $4$ ballots of the form $C\succ A$. Let $\mathcal{B}_{2}$ consist of the sub-election with 3m103m-10 ballots: m2m-2 ballots of the form $A\succ B$; m2m-2 ballots of the form $B\succ A$; m/24m/2-4 ballots of the form $C\succ A$; and m/22m/2-2 ballots of the form $C\succ B$. Then $\mathcal{B}_{1}$ and $\mathcal{B}_{2}$ form a partition of $\mathcal{B}$, and candidate $B$ wins both of these sub-elections. Since $m$ can be arbitrarily large, this implies there exists an election demonstrating a reinforcement paradox in which one of the sub-elections consists of an arbitrarily small fraction of the ballots in the original election.
@@ 4. Results under various models
In this section we use different models of voter behavior to generate millions of synthetic elections. Among the many possible models we could use, we choose models which have been used extensively in prior social choice research. For each model, we generate 100,000 profiles and check whether a generated election is susceptible to a reinforcement paradox. Thus, our primary tool for providing (estimated) probabilities is Monte Carlo simulation, although we can provide exact probabilities using other means for the impartial anonymous culture (IAC) model.
We are interested in probabilities for relatively large electorates, as the real-world elections we analyze in the next section have electorate sizes ranging from a few hundred to almost one million (although most of the elections contain 20,000 or fewer voters). Thus, our simulations use an electorate size of $V=1001$ voters. We also ran simulations with $V\in\{601,2003,3901\}$, but the estimated probabilities were not markedly different for these other choices of $V$. With the exception of the IAC model, in this section we consider probabilities only for the complete ballot case.
@@ 4.1. Descriptions of the Models
We use two different types of models: one-dimensional spatial models and models which sample over the ballot simplex. In the spatial framework, we consider three different models. The first model, which we denote 1D Spatial, places three candidates and $V=1001$ voters on the real line using the standard normal distribution, and a voter’s preference ranking is determined by their Euclidean distance from each candidate. For example, if the three candidates $A$, $B$, and $C$ are positioned by the normal distribution at 1.2-1.2, 0.1-0.1, and 1.1 respectively, then a voter positioned at 0.3 would cast a ballot with $B$ ranked first, $C$ ranked second, and $A$ ranked third. Such one-dimensional spatial models have been widely used in the social choice literature [17, 18, 29, 51].
Because of the increased political polarization in the United States and elsewhere, we also use two bimodal spatial models to investigate the potential effects of polarization on the probability of the reinforcement paradox. Under the second spatial model, which we denote 1D Bimodal, we create a bimodal distribution from an equally weighted combination of the normal distribution centered at 2-2 and the normal distribution centered at 2, both with standard deviation 1. See Figure 1 (top) for a visualization of this bimodal distribution. We then place candidates and voters along the line using the generated (normalized) bimodal distribution as a probability distribution, similar to 1D Spatial. Under the third spatial model, which we denote 1D Bimodal Weighted, we use the same two normal distributions to generate the bimodal distribution but give 50% more weight to the left distribution, resulting in a bimodal distribution as shown Figure 1(bottom). We use this model to capture polarization when one side of the political spectrum is stronger than the other. (We do not investigate the effect of changing the variance.)
@@
[…]
@@ 4.2. Results
We begin by providing exact probabilities under Dir(1)(1), or IAC model. Because of its simplicity and prevalence in the social choice literature, researchers have dedicated significant effort to developing tools for calculating probabilities under this model. One of the fruits of these labors is the software package Normaliz [9], which can provide limiting probabilities under IAC in three-candidate elections given a set of linear constraints. In the complete ballot case, Normaliz determines the proportion of integer values of $X_{1},X_{2},Y_{1},Y_{2},Z_{1}$ and $Z_{2}$ that satisfy the inequalities in Proposition 1 given a fixed number of voters $V$, as $V\rightarrow\infty$. In the partial ballot case, Normaliz, does the same thing for $X_{1},X_{2},X_{3},Y_{1},Y_{2},Y_{3},Z_{1},Z_{2}$ and $Z_{3}$ that satisfy the inequalities in Proposition 2.
The probabilities in the following proposition are thus exact (rather than approximated through simulation) and are obtained directly from Normaliz by inputting the appropriate sets of inequalities. The conditional probabilities are obtained by adding an additional requirement to limit the number of elections to those without a majority candidate.
@@
The limiting probabilities that an election is susceptible to a reinforcement paradox under the Dir(1)(1) model (the IAC model) are as shown in Table 11 for complete and partial ballots. The “Cond. Prob.” probabilities are conditioned on the absence of a majority candidate.
@@
Table 11. Probabilities under the Dir(1)(1) (IAC) model. The top two rows correspond to the complete ballot case; the bottom two rows correspond to the partial ballot case.
The probabilities of one or more paradoxes occurring are similar in the complete and partial ballots cases. When conditioned on the absence of a majority candidate, the probabilities are significantly lower in the partial ballot case. This is because the probability an election does not contain a majority candidate is $145/256=56.64\%$ in the 8-simplex but is only $7/16=43.75\%$ in the 5-simplex. Overall, the probabilities are fairly high. This is largely because they indicate the percentage of elections in which there exists a partition into two sub-elections demonstrating a reinforcement paradox; the numbers do not indicate anything about the number of such partitions in any given election.
[…]
@@
Interestingly, the corresponding probabilities under 1D Spatial and 1D Bimodal are approximately equal, perhaps because both models are symmetric about zero. When we unbalance the distribution under 1D Bimodal Weighted, the probabilities change significantly. In particular, 1D Bimodal Weighted is the only model in this study under which it is more likely to observe a plurality loser paradox than a runner-up paradox. We do not know with certainty the reason for this, but we speculate that it is due to different responses to the so-called “center squeeze” [49]. In one-dimensional three-candidate spatial models, the central candidate will often be the plurality loser even if they are closest to the median voter. Using the notation of Section 2, if $C$ is the central candidate then $X_{1}=Y_{1}=0$ since $C$ must be second-ranked by all voters. Hence the conditions in Case (1) of Proposition (1) cannot be satisfied (since they imply $B$ has the majority of the first-place vote). The conditions in Case (2) reduce to
@@
If the distribution of voters is symmetric, $Z_{1}$ and $Z_{2}$ are likely to be similarly-sized, and so the rightmost inequality will be satisfied whenever $Y>X$. Thus, a runner-up (RU) paradox will occur often when the IRV and plurality winners are distinct.
If the distribution of voters is not symmetric, as with the 1D Bimodal Weighted model, the left and right-wing candidates are likely to be near the left and right peaks of the distribution respectively, with the centrist in between but on the “upward slope” of the larger peak (the left one in Figure 1). Thus, the quantity $Z_{1}-Z_{2}$ is more likely to be larger than in the symmetric case, decreasing the likelihood of Case (2) occurring. A similar argument holds if $A$ or $B$ is the central candidate.
For the plurality loser (PL) paradox to occur, it is easily seen that $C$ must be the central candidate. In this case, $C$ is again likely to be on the upward slope of the larger peak in the 1D Bimodal Weighted model (and hence $Z$ is likely to be larger) than in the 1D Bimodal model. So the probability of the plurality loser (PL) paradox occurring is greater in the nonsymmetric case than the symmetric.